40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 212 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  6318   Fri Feb 24 19:25:43 2012 jamieUpdateLSCALS X-arm beatbox added, DAQ channels wiring normalized

I have hooked the ALS beatbox into the c1ioo DAQ.  In the process, I did some rewiring so that the channel mapping corresponds to what is in the c1gcv model.

The Y-arm beat PD is going through the old proto-DFD setup.  The non-existant X-arm beat PD will use the beatbox alpha.

Y coarse I (proto-DFD) --> c1ioo ADC1 14 --> C1:ALS_BEATY_COARSE_I
Y fine   I (proto-DFD) --> c1ioo ADC1 15 --> C1:ALS_BEATY_FINE_I
X coarse I (bbox alpha)--> c1ioo ADC1 02 --> C1:ALS_BEATX_COARSE_I
X fine   I (bbox alpha)--> c1ioo ADC1 03 --> C1:ALS_BEATX_FINE_I

This remapping required coping some filters into the BEATY_{COARSE,FINE} filter bank.  I think I got it all copied over correctly, but I might have messed something up.  BE AWARE.

We still need to run a proper cable from the X-arm beat PD to the beatbox.

I still need to do a full noise/response characterization of the beatbox (hopefully this weekend).

  6317   Fri Feb 24 19:18:28 2012 kiwamuUpdateLSCY arm + PRMI : how they should look like

I calculated how the DC signals should look like in the Y arm PRMI configuration.

The expected signals are overlaid in the same plot as that of shown in #6313.

You can see there are disagreements between the observed and expected signals in the plot below at around the time when the arm is brought to the resonance.

 

(expected behaviors)

  •  TRY: At the end it should be at 1 (remember TRY is normarlized) and should not go more than that, since the power-recycling is in a weird situation and it is not fully recycling the power.
  •  ASDC: It should become brighter at the end because the arm cavity flips the sign of the reflected light and hence the dark port must be on a bright fringe.
  •  REFLDC: It will decrease a little bit because the arm cavity and MICH try to suck some amount of the power into the interferometer.

 

expected_time_series.png

Quote from #6313

The figure below shows the time series of the Y arm + PRMI trail.

  6316   Fri Feb 24 18:59:04 2012 JenneUpdateComputersPyNDS and a Plot

Quote:

Power Spectral Density plot using PyNDS, comparing 5 fast data channels for ETMX.

 Is there any stuff to install, etc?  Y'know, for those of use who don't really know how to use computers and stuff....

  6315   Fri Feb 24 18:37:13 2012 ranaUpdateLSCY arm + PRMI part II

Quote:

I tried the Yarm + PRMI configuration again.

The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.

 

 Isn't the point that the 11 and 55 MHz signals have the carrier effect, but the 3f signals are better?

  6314   Fri Feb 24 16:10:48 2012 mikeUpdateComputersPyNDS and a Plot

Power Spectral Density plot using PyNDS, comparing 5 fast data channels for ETMX.

**EDIT** Script here:

import nds
import numpy as np
import matplotlib.pyplot as plt
import time
daq=nds.daq('fb', 8088)
channels=daq.recv_channel_list()
e=0
start=int(time.time()-315964819)
rqst=['C1:SUS-ETMX_SENSOR_UR','C1:SUS-ETMX_SENSOR_UL','C1:SUS-ETMX_SENSOR_LL','C1:SUS-ETMX_SENSOR_LR','C1:SUS-ETMX_SENSOR_SIDE']    #Requested Channels
for c in channels:
    if c.name in rqst:
        daq=nds.daq('fb', 8088)
        data=daq.fetch(start-100, start, c.name)
        vars()['psddata'+str(e)], vars()['psdfreq'+str(e)]=plt.psd(data[0],NFFT=16384,Fs=c.rate)
        vars()['label'+str(e)]=c.name
        e+=1
plt.figure(1)
plt.clf()
plt.title('PSD Comparison')
plt.grid(True, which='majorminor')
plt.xlabel(r'Frequency $Hz$')
plt.ylabel(r'Decibels $\frac{dB}{Hz}$')       
for x in np.arange(0,e):
    plt.loglog(psdfreq0, 10*vars()['psddata'+str(x)], label=vars()['label'+str(x)])
plt.legend()
plt.show()

Attachment 1: PSD_Comparison.png
PSD_Comparison.png
  6313   Fri Feb 24 15:01:31 2012 kiwamuUpdateLSCY arm + PRMI part II

The figure below shows the time series of the Y arm + PRMI trail.

time_series.png

(Top plot )

  Normalized TRY (intracavity power). It is normalized such that it shows 1 when the arm is locked with the recycling mirrors misaligned.

(Middle plot)

ASDC and REFLDC in arbitrary unit.

(Bottom plot)

The amount of the arm length detuning observed at the fine frequency discriminator.

 

(Sequence)

At t = 20 sec, the amount of detuning was adjusted so that the cavity power goes to the maximum. At this point the PRM was misaligned.

At t = 30 sec, the cavity length started being slowly detuned to 10 nm. As it is being detuned the intracavity power goes down to almost zero.

At t = 45 sec, the alignment of PRM was restored. Because of that, the REFLDC and ASDC diodes started receiving a large amount of light.

At t = 85 sec, the PRCL and MICH were locked. The REFLDC signal became a high value as the carrier light is mostly reflected. The ASDC goes to a low value as the MICH is kept in the dark condition.

At t = 100 sec, the length started being slowly back to the resonance while the PRMI lock was maintained.

At t = 150 sec, the lock of the PRCL and MICH were destroyed. With the arm fully resonance, I wasn't able to recover the PRMI lock with the same demod signals.

Quote from #6310

I tried the Yarm + PRMI configuration again.

The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.

I will post the time series data later.

 

  6312   Fri Feb 24 08:06:52 2012 steveUpdateSUSsus restored

Quote:
The following optics were kicked:
MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS
Fri Feb 24 04:11:15 PST 2012
1014120690
 
Steve (or anyone), can you restore the watchdogs when you come to the lab in the morning ?

Quote from #6305

Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.

 All suspentions were restored and MC locked. PRM side osem  RMS motion was high.

Atm2, Why the PRM is 2x as noisy as the SRM ?

Attachment 1: freePRM.png
freePRM.png
Attachment 2: noisyPRM.png
noisyPRM.png
  6311   Fri Feb 24 04:12:44 2012 kiwamuUpdateSUSfreeswing test
The following optics were kicked:
MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS
Fri Feb 24 04:11:15 PST 2012
1014120690
 
Steve (or anyone), can you restore the watchdogs when you come to the lab in the morning ?

Quote from #6305

Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.

  6310   Fri Feb 24 03:58:13 2012 kiwamuUpdateLSCY arm + PRMI part II

I tried the Yarm + PRMI configuration again.

The PRMI part was locked, but it didn't stay locked during the Y arm was brought to the resonance point.

I will post the time series data later.

 

(locking of the PRMI part)

Tonight I was able lock the PRMI when the arm was off from the resonance by 10 nm (#6306).

This time I used REFL11Q to lock the MICH instead of the usual AS55Q because the MICH didn't stay locked with AS55Q for some reason.

The PRCL was held by REFL33I as usual.

Also I disabled the power normalization for the error signals because it could do something bad during the Y arm is borough to the resonance.

In order to reduce the number of the glitches, PRM was slightly misaligned because I knew that the lower finesse gives fewer glitches.

  6309   Thu Feb 23 10:53:54 2012 steveUpdateSUS oplev PIT transfer functions

 Oplev transfer functions PIT UGF were optimized to be at 2-3 Hz with 60 degree minimum phase margin by adjuting oplev gains.

Additional Notes by KI:

  • The PRM oplev has a tailored 3.3 Hz resonant gain in order to calm down a wobble during the lock acquisitions.
  • Also in the PRM oplev a 35 Hz elliptic cut-off filter wasn't  activated at the time when Steve measured it.
  • In both ITMs, elliptic cut-off filters seem to have higher corner frequencies compered with the others.
    • I guess these settings are from the old days.
  • ETMs and ITMs have whitening filters while the rest of the suspensions don't.
    • Without the whitening filters, normally the signals above 30 Hz are covered by some electrical noises or perhaps He-Ne laser intensity noise (#5630).
      • This is why we usually use the 35 Hz elliptic filters to roll off the control noises.
    • Since the ETMs and ITMs have whitening filters they potentially can have slightly higher corner frequencies in the elliptic filters.
      • Of course the corner frequencies need to be re-designed in terms of the amount of noise injection to the longitudinal motion.

 

Attachment 1: oplPITtrans.pdf
oplPITtrans.pdf
  6308   Thu Feb 23 09:09:33 2012 steveUpdateSAFETYsafety checks

Emergency exit lights were inspected:  2 out of 13 batteries have to be replaced

One of the Halon fire extinguishers needs to be recharged out of 8

Please do participate in preparation for the upcoming safety audit on Feb 28

  6307   Thu Feb 23 02:20:07 2012 ZachUpdateSUSwacky state of SUS input matrices

This reminds me that the whole Dr. SUS situation never got taken care of. Where I left off, I was having issues pulling 40m data with NDS2 (which is what all the diagonalization scripts use).

What is the deal with 40m+NDS2? If it is till no-go, can we have a consensus on whether this is too important to wait for? If so, I will rewrite the scripts to use NDS and we can upgrade to NDS2 once we can prove we know how to use it.

 

Quote:

While Kiwamu and I were trying to investigate the the vertex glitches we were noticing excess noise in ITMX, which Kiwamu blamed on some sort of bad diagonalization.  Sure enough, the ITMX input matrix is in the default state [0], not a properly diagonalized state.  Looking through the rest of the suspensions, I found PRM also in the default state, not diagonalized.

We should do another round of suspension diagonalization.

Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.

[0]

0.25 0.25 0.25 0.25 0
1.66 1.66 -1.66 1.66 0
1.66 -1.66 -1.66 1.66 0
0 0 0 0 1

 

  6306   Wed Feb 22 19:45:33 2012 kiwamuUpdateLSChow much length offset do we need ?

I did a quick calculation to see if the offset of the arm length which I tried last night was reasonable or not.

The conclusion is that the 20 nm offset that i tried could be a bit too close to a resonance of the 55 MHz sidebands.

A reasonable offset can be more like 10 nm or so where the phases of all the laser fields don't get extra phases of more than ~ 5 deg.


 

The attached plot shows where the resonances are for each sideband as a function of the displacement from the carrier's resonance.

armresonance.png

The red solid line represent the carrier, the other solid lines are for the upper sidebands and the dashed lines are for the lower sidebands.

The top plot shows the cavity power and the bottom plot shows how much phase shift the fields get by being reflected by the arm cavity.

Apparently the closest resonances to the the main carrier one are that of the 55 MHz sidebands, and they are at +/- 22 nm.

So if we displace the arm length by 22 nm, either of the 55 MHz sidebands will enter in the arm cavity and screw up the sensing matrix for the 55 MHz family.

Quote from #6304

In all the configurations I displaced the Y arm by 20 nm from the resonance.

  6305   Wed Feb 22 16:55:16 2012 JamieUpdateSUSwacky state of SUS input matrices

While Kiwamu and I were trying to investigate the the vertex glitches we were noticing excess noise in ITMX, which Kiwamu blamed on some sort of bad diagonalization.  Sure enough, the ITMX input matrix is in the default state [0], not a properly diagonalized state.  Looking through the rest of the suspensions, I found PRM also in the default state, not diagonalized.

We should do another round of suspension diagonalization.

Kiwamu (or whoever is here last tonight): please run the free-swing/kick script (/opt/rtcds/caltech/c1/scripts/SUS/freeswing) before you leave, and I'll check the matrices and update the suspensions tomorrow morning.

[0]

0.25 0.25 0.25 0.25 0
1.66 1.66 -1.66 1.66 0
1.66 -1.66 -1.66 1.66 0
0 0 0 0 1

  6304   Wed Feb 22 13:28:22 2012 kiwamuUpdateLSCY arm + central part locking

Last night I tried the "Y arm + central part" locking again. Three different configuration were investigated :

  •  Y arm + DRMI
  •  Y arm + PRMI
  •  Y arm + MICH

In all the configurations I displaced the Y arm by 20 nm from the resonance.

As for the DRMI and PRMI configurations I wasn't able to acquire the locks.

As for the MICH configuration, the MICH could be locked with AS55. But after bringing the Y arm to the resonance point the lock of MICH was destroyed.

  6303   Wed Feb 22 01:53:57 2012 kiwamuUpdateLSCupdate on glitch table

I tried SRMI. The glitch rate wasn't as high as that of PRMI but it happened once per 10 sec or so.

 

 

 Yarm

(POY11 -->

ETMY)

Xarm

(POX11 --> ETMX)

MICH

(AS55-->BS)

or

(AS55 --> ITMs)

Half PRMI

(REFL11 --> PRM)

or

(REFL33 --> PRM)

low finesse PRMI

(ASDC --> ITMs)

(REFL33 --> PRM)

PRMI (carrier)

(AS55 --> ITMs)

(REFL33 --> PRM)

PRMI (sideband)

(AS55 --> ITMs)

(REFL33 --> PRM)

SRMI(NEW)

(AS55-->ITMs)

(REFL11I --> SRM)

DRMI
AS55 NO NO NO NO glitch (depends on finesse)
glitch glitch glitch glitch
REFL11 NO NO NO NO glitch (depends on finesse)
glitch glitch glitch glitch
REFL33 NO NO NO NO - glitch glitch glitch glitch
REFL55 NO NO NO NO glitch(depends on finesse) glitch glitch glitch glitch
REFL165 NO NO NO - - - - - -
POX11 - NO NO NO  - glitch glitch - glitch
POY11 NO - NO NO  - glitch glitch - glitch
POP55 - - - -  -  - -   -
                   

 

Quote from #6284

I updated the table which I posted some time ago (#6231). The latest table is shown below.

It seems that the glitches show up only when multiple DOFs are locked.

 

  6302   Tue Feb 21 22:06:18 2012 jamieUpdateLSCbeatbox DFD installed in 1X2 rack

I have installed a proto version of the ALS beatbox delay-line frequency discriminator (DFD, formally known as MFD), in the 1X2 rack in the empty space above the RF generation box.

That empty space above the RF generation box had been intentionally left empty to provide needed ventilation airflow for the RF box, since it tends to get pretty hot.  I left 1U of space between the RF box and the beatbox, and so far the situation seems ok, ie. the RF box is not cooking the beatbox.  This is only a temporary arrangement, though, and we should be able to clean up the rack considerably once the beatbox is fully working.

For power I connected the beatbox to the two unused +/- 18 V Sorensen supplies in the OMC power rack next to the SP table.  I disconnected the OMC cable that was connected to those supplies originally.  Again, this is probably just temporary.

Right now the beatbox isn't fully functioning, but it should be enough to use for lock acquisition studies.  The beatbox is intended to have two multi-channel DFDs, one for each arm, each with coarse and fine outputs.  What's installed only has one DFD, but with both coarse and fine outputs.  It is also intended to have differential DAQ outputs for the mixer IF outputs, which are not installed in this version.

The intended design was also supposed to use a comparator in the initial amplification stages before the delay outputs.  The comparator was removed, though, since it was too slow and was limiting the bandwidth in the coarse channel.  I'll post an updated schematic tomorrow.

I made some initial noise measurements:  with a 21 MHz input, which corrseponds to a zero crossing for a minimal delay, the I output is at ~200 nVrms/\sqrt{Hz} at 5 Hz, falling down to ~30 nVrms about 100 Hz, after which it's mostly flat.  I'll make calibrated plots for all channels tomorrow.

The actual needed delay lines are installed/hooked up either.  Either Kiwamu will hook something up tonight, or I'll do it tomorrow.

  6301   Tue Feb 21 18:39:11 2012 kiwamuUpdateGreen LockingNew BBPDs installed

Two new BBPDs have been installed on the PSL table.

The first one was installed by Koji a few days ago, and I stalled the second one today.

They will serve as beat-note detectors for the green locking.

Next step : I have to lay down a long SMA cable which goes from the BBPD to the IOO rack.

  6300   Tue Feb 21 16:10:29 2012 kiwamuUpdateIOOdegradation in input PZT1

PZT1, the one with Koji's custom mid-HV driver (#5447), is getting degraded.

The movable range in the pitch direction became narrower than what it used to be (maybe a factor of 3 estimated by looking at the beam spots).

I think we should raise the priority level of the active TTs for the next vent.

 

I have been having a feeling that the PZT1 response is getting smaller since the end of the last year, but now I am confident

because I could see the difference between the movable ranges of Yaw and Pitch, and they used to have approximately the same amount of the movable ranges.

Right now this is not a serious issue as the beam pointing determined by the MC alignment is so good that the Pitch range doesn't rail.

I won't be surprised if it becomes completely immovable in 3 month.

  6299   Tue Feb 21 08:33:16 2012 steveUpdateGreen Lockingperiscope adapter plate

Two extender plates ready for cleaning. The existing optical table tops have 38" OD. Using two of these the OD will be 44"

Attachment 1: percplate.PDF
percplate.PDF
  6298   Tue Feb 21 04:30:02 2012 kiwamuUpdateLSCY arm + PRMI

I tried the "Yarm + PRMI" configuration to see what happens.
The Y arm was locked at a resonance and held with the ALS technique.
On the other hand, the X arm was freely swinging.

I briefly tried severl demod signals to calm down the central part, but didn't succeed.
Now I feel I really want to have the X arm locked with the ALS technique too.
Give me the beat-box !

The attached screen shot shows the transmitted light of both arms as a function of time.
TRY is always above 1, since it was kept at a resonance.
Sometimes TRY went to 50 or so.

Untitled.png

  6297   Sat Feb 18 18:29:38 2012 DenUpdateAdaptive Filteringonline filtering

I tried to filter MC_F from seismic noise measured by GUR1 seismometer. I've used 8000 tap filter, downsample ratio=8, delay=1. In the Figure the output of the filter is presented with MC_F signal.

output.pdf

We can see that output is close to the MC_F, but the phase for some reason is not zero. It should not be at 1 Hz - 10 Hz due to the actuator. But below these frequencies I do not see any reasons for the output phase to differ from MC_F phase. But it is possible, the phase of the actuator is evaluated very rough and the adaptive filter can't match it.

  6296   Sat Feb 18 17:01:26 2012 DenUpdateAdaptive Filteringstatic variables

In order to prevent different DOF from redetermining static variables in the adaptive code, I've created a separate code for each DOF with the name ADAPT_XFCODE_{$DOF}.c

I've provided the links for these files in the c1oaf.mdl, compiled and run it. Now there are no conflicts between DOFs.

  6295   Sat Feb 18 16:58:59 2012 kiwamuUpdateIOOMC suspension realigned

[ Den / Kiwamu]

 We have realigned the MC suspensions so that the WFS servos are smoothly engaged.

Now it seems working fine. The beam pointing to the interferometer also looks okay.

The WFSs control kept failing to engage the servos because of large misalignments in the MC suspensions.

When the TEM00 was locked, the transmitted light was only about 1200 counts and the reflected light was about 2.8 counts.

We tweaked MC1, MC2 and MC3.

Quote from #6294

When I came to the 40m this afternoon, the MC was unlocked. Here is the trend of MC_F for last 2 hours

  6294   Sat Feb 18 12:23:09 2012 DenUpdateIOOMC

When I came to the 40m this afternoon, the MC was unlocked. Here is the trend of MC_F for last 2 hours

mc_trend.png

C1:PSL-PMC_PMCTRANSPD = 0.800

Should I just disable the auto locker or try to realign it?

  6293   Fri Feb 17 04:45:48 2012 kiwamuUpdateLSCsensing matrix of PRMI

I locked the PRMI with the AS55I and Q combination.

It seems the glitche rate decreased,

but I am not 100 % sure because the rest of the demod signals (i.e. REFL11 and etc) were showing relatively big signals (noise ?), which may cover the glitches.

Also the optical gain of PRCL at AS55I doesn't agree with my expectation based on the obtained sensing matrix (#6283).

It looks too low and lower than the measured sensing matrix by a factor of 50 or so.

I will continue working on this configuration tomorrow and then move on to the SRMI locking as a part of the glitch hunting activity.

Quote from #6287

So why don't you use AS55I and Q for the control of PRMI???

  6292   Fri Feb 17 01:02:22 2012 kiwamuUpdateIOOMC is back to normal

[Koji / Kiwamu]

 The MC is now back to normal. The beam pointing to the interferometer is good.

There were two different issues :

  • A mechanical mount was in the MC WFS path.
  • There were some loose connections in the SUS rack

 

Slid have we the position of the mechanical mount. Nicely the WFS beam go through now.

And also I pushed all the connectors associated with the MC SUS OSEMs in the SUS rack.

After pushing the connectors, the MC1 OSEM readouts dramatically changed, which actually more confused us.

As shown in the 3 hours trend below, the OSEM readouts have changed a lot (shown in the middle of the plot with arrows). Some bumps after the steps correspond to our alignment efforts.

MCconnections.png

Quote from #6291

The MC became crazy again.

 

  6291   Thu Feb 16 23:12:55 2012 kiwamuUpdateIOOMC unlocking frequently

The MC became crazy again.

It seems that there were corresponding steps in the OSEM signals. Look at the one-day trend posted below.

MCsteps.png

  6290   Thu Feb 16 21:13:07 2012 SureshUpdateElectronicsREFL165 repair: PD replaced, DC response checked with a torch light

[Koji, Suresh]

Kiwamu mentioned that REFL165 is not responding and its DC out seems saturated at 9V.  Koji and I checked to see if changing the power supply to the PD changed its behaviour. It did not.  

I then look a close look at the PD and found that the front window of the PD was not clear and transparent.  There was a liquid condensation inside the window, indicating an over heating of the PD at some point.  It could have arisen due to excessive incident power.  The pic below shows this condensation:

PC_30641_old.jpg

 

I also checked the current flowing through the reverse bias voltage line.  There was a voltage drop of 3V across R22 (DCC D980454-01-C)   indicating a 150mA of current through the PD.  This is way too much above the operating current of about 20mA.   The diode must have over heated.

I pulled out the old PD out and installed a new one from stock.  The pic below shows the clear window of a new PD.

PD_30641_new.jpg

After changing the PD I checked the DC output voltage while shining a torch light on to the PD.  It showed an output of about 30 to 40 mV.  This seemed okay because the larger 2mm photodiodes showed ~100mA DC output with the same torch.Below is the current state of the ckt board.

IMG_0548.JPG

 

I will tune the PD to 165 MHz tomorrow and measure its transimpedance.

  6289   Thu Feb 16 13:12:30 2012 ranaUpdateLSCsensing matrix of PRMI

Quote:

I think I have told a lie in the last meeting -- the measured sensing matrix doesn't look similar to what Optickle predicts.

Smells like something is very wrong.

 Those Radar plots are awesome. Even more awesome would be if they were in units of W/m (so that it can be directly compared with Optickle) and so that the numbers are useful even 1 year from now. Otherwise, we will lose the RF transimpedance information and thereby lose everything.

Also, please post the provenance of the counts->V calibration.

  6288   Thu Feb 16 09:59:16 2012 steveUpdateASCIP- ANG

Initial pointing or IP-ANG is a pointing monitor of the MC. This beam is launched after the second  pzt  steering mirror.

IP-ANG  is missing the pick up mirror by a few inches at ETMYchamber

1000 days plot show last appearance in Feb 2010

Attachment 1: lastIPang.png
lastIPang.png
  6287   Thu Feb 16 07:38:24 2012 KojiUpdateLSCsensing matrix of PRMI

So why don't you use AS55I and Q for the control of PRMI???

  6286   Thu Feb 16 04:29:30 2012 kiwamuUpdateLSCupconversion noise from BS motion

Sometimes ago I reported that there have been a kind of upconversion noise when PRM was excited (#6211).

This time I found another one, which showed up when BS was excited.

Assuming this is related to some kind of scattering process and also assuming this is from the same scattering body as that for the PRM driven case,

we may be able to localize and perhaps identify the scattering body.

 

 

(Measurement Condition)

All the suspended optics are intentionally misaligned except for ITMY so that the laser directly goes through to the dark port without any interference.
Then BS_POS is excited at 3 Hz with amplitude of 1000 counts by an oscillator in the realtime lockin system.
 I also excited PITCH and YAW of BS and found that driving the angular motions didn't produce any upconversion noise.
 I didn't excite ITMY to do the same test because I was too lazy.

(Noise spectrum)

The plot below shows the upconversion noise observed at AS55 and REFL11.
The reference curves were obtained when no excitation were applied on BS_POS.
It is obvious that the AS55 signal shows a typical upconversion behavior.
 
Untitled.png

 

  6285   Thu Feb 16 04:02:16 2012 kiwamuUpdateLSCinsane REFL165 DC output

I found that the DC monitor of the REFL165 was showing 9 V regardless of how much laser power goes to the diode.

I am worried about whether the RF output is also broken.

It needs to be checked and I will leave this to Suresh as one of his morning tasks.

  6284   Thu Feb 16 03:47:16 2012 kiwamuUpdateLSCglitch table

I updated the table which I posted some time ago (#6231). The latest table is shown below.

It seems that the glitches show up only when multiple DOFs are locked.

Interesting thing is that when the low finesse PRMI is locked with a big MICH offset (corresponding to a very low finesse) it doesn't show the glitches.

Qualitatively speaking, the glitch rate becomes higher as the finesse increases.

I will try SRMI tomorrow as this is the last one which I haven't checked the presence of the glitches.

 

 

 Yarm

(POY11 -->

ETMY)

Xarm

(POX11 --> ETMX)

MICH

(AS55-->BS)

or

(AS55 --> ITMs)

Half PRMI

(REFL11 --> PRM)

or

(REFL33 --> PRM)

low finesse PRMI

(ASDC --> ITMs)

(REFL33 --> PRM)

PRMI (carrier)

(AS55 --> ITMs)

(REFL33 --> PRM)

PRMI (sideband)

(AS55 --> ITMs)

(REFL33 --> PRM)

DRMI
AS55 NO NO NO NO glitch (depends on finesse)
glitch glitch glitch
REFL11 NO NO NO NO glitch (depends on finesse)
glitch glitch glitch
REFL33 NO NO NO NO - glitch glitch glitch
REFL55 NO NO NO NO glitch(depends on finesse) glitch glitch glitch
REFL165 NO NO NO - - - - -
POX11 - NO NO NO  - glitch glitch glitch
POY11 NO - NO NO  - glitch glitch glitch
POP55 - - - -  -  - - -
                 

 

 

  6283   Wed Feb 15 17:15:33 2012 kiwamuUpdateLSCsensing matrix of PRMI

I think I have told a lie in the last meeting -- the measured sensing matrix doesn't look similar to what Optickle predicts.

Smells like something is very wrong.

 

Measured sensing matrix

The measured matrix are shown in the diagram below.
The lengths of arrows corresponds to the signal strength in unit of V/m. The radial axis in in log scale.
The angle of arrows corresponds to their best demodulation phases.
sensingMAT.png

        Some obvious things:

  •  REFL11 : The separation angle between MICH and PRCL is narrow and it is far from the ideal 90 degree. This doesn't agree with the simulation.
  •  REFL33:  The MICH and PRCL signals are almost degenerated in their demodulation phase.
  •  REFL55 :  It shows non-90 degree separation. This doesn't agree with the simulation.
  •  REFL165 : The separation is close to 90 degree, but the signals are small. And I am not sure if the MICH signal is real or just noise.
  •  AS55 : Somehow it shows a nice 90 degree separation, but this result doesn't agree with the simulation.

 


Expected sensing matrix from a simulation

For a comparison here is a result from an Optickle simulation.
This time the radial unit is W/m instead of V/m, but they are qualitatively the same unit.
The radial axis is in log, so when it says 2, it means 10^2 [W/m].
PRMI_Optickle.png
 
 Simulation setup:
l_PRC  = 6.760 (see #4064)
l_asy  = 0.0364  (see #4821)
loss per optic = 50 ppm
 

Measurement

  •  Locked PRMI with the carrier anti-resonating in PRCL.
  •  Adjusted the control gains for both the MICH and PRCL control to have UGFs at ~ 100 Hz.
  •  Put a 30 dB notch filter  in each control servo at 283.1 Hz where an excitation signal will be.
  •  Excited PRCL and MICH at different time via the realtime lockng in the LSC front end. The amplitude is 1000 counts and the frequency is at 238.1 Hz.
    • For the MICH excitation, I have coherently and differentially excited ITMs
  •  Used DTT to take a transfer function (transfer coefficients at 283.1 Hz) from the lockin oscillator to each LSC demodulated signal.
    • Including AS55I/Q, REFL11I/Q, REFL33I/Q, REFL55I/Q and REFL165I/Q.
  •  Calibrated the obtained transfer functions from unit of counts/counts to V/m using the actuator response (#5637)

Quote from #6281
I have measured the sensing matrix of PRMI.
It seems that the MICH signal in the 3f ports (REFL33 and REFL165) were quite tiny, and because of that it is very tough to use them for the actual MICH control.
The data is coming soon.

  6282   Wed Feb 15 11:34:01 2012 steveUpdate under the shouth end optical table

I added an U channel based bottom shelf at the south end today.

Attachment 1: P1080535.JPG
P1080535.JPG
Attachment 2: P1080537.JPG
P1080537.JPG
  6281   Wed Feb 15 05:29:22 2012 kiwamuUpdateLSCsensing matrix of PRMI

I have measured the sensing matrix of PRMI.

It seems that the MICH signal in the 3f ports (REFL33 and REFL165) were quite tiny, and because of that it is very tough to use them for the actual MICH control.

The data is coming soon.

  6280   Tue Feb 14 17:09:05 2012 steveUpdateGeneralreflectivity of green welding glass

 Schott, Athermal green welding glass, shade #14 reflectivity was measured in 1.2W,  ~1 mm diameter  beam of MC reflected.

The P polarization measurement  was done with the help of half wave plate and PBC

 

 

Attachment 1: gg#14ref.pdf
gg#14ref.pdf
Attachment 2: gg14refl1Wps.png
gg14refl1Wps.png
  6279   Tue Feb 14 15:52:11 2012 JenneUpdateAuxiliary lockingYarm fiber returned to ATF

[Frank, Jenne]

We extracted the fiber that Suresh and Sonali laid over the summer, for the IR beat for the Ygreen laser, and Frank took it back to Bridge to be used in the new fiber distributed reference laser setup.

  6278   Tue Feb 14 08:22:27 2012 steveUpdateSUSsus damping restored

ITMX, PRM and BS watchdogs are tripped. They were restored.

Stable MC was disabled so I can use MC_REFL 1 W beam to measure green glass .

  6277   Mon Feb 13 12:02:17 2012 KojiUpdatePEMseismic BLRMS loud too

I reported the procedure to add slow channels to the FB. I guess you already have done Step.1

http://nodus.ligo.caltech.edu:8080/40m/5991

Quote:

 Tragically, this is more tricksy than I would have thought. The channels we need are "cdsEpicsOutput"s in the model.  They don't show up in Dataviewer (fast or slow channels) or the regular fast channel .ini file.  Jamie and I don't remember where these channels live and how to get them saved to frames.  I'm on top of it though.

 

  6276   Mon Feb 13 11:30:51 2012 JenneUpdatePEMseismic BLRMS loud too

Quote:

Quote:

 

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

 Seems like a problem to solve on Monday so that we don't end up without trends like this again.

 Tragically, this is more tricksy than I would have thought. The channels we need are "cdsEpicsOutput"s in the model.  They don't show up in Dataviewer (fast or slow channels) or the regular fast channel .ini file.  Jamie and I don't remember where these channels live and how to get them saved to frames.  I'm on top of it though.

I did notice however, that the striptool for seismic trends is showing the wrong channels for 3-10 and 10-30 Hz.  The other 3 channels are correctly the output after the sqrt is taken, but those two (orange and red on striptool) are before the sqrt, but after the bandpass and low pass.  I'll fix that now...

  6275   Fri Feb 10 23:58:30 2012 ranaUpdatePEMseismic BLRMS loud too

Quote:

 

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

 Seems like a problem to solve on Monday so that we don't end up without trends like this again.

  6274   Fri Feb 10 23:19:09 2012 kiwamuUpdateIOOcross talk causing fake seimometer signals

[ Koji / Kiwamu ]

The frequent unlock of the MC are most likely unrelated to ground motion.

Although the reason why MC became unstable is still unclear.

 

 

There are two facts which suggest that the ground motion and the MC unlock are unrelated :

(1) It turned out that the seismometer signals (C1:PEM-SEIS-STS_AAA ) have a big cross talk with the MC locking signals.

    For example, when we intentionally unlocked the MC, the seismometer simultaneously showed a step-shaped signals, which looked quite similar to what we have observed.

    I guess there could be some kind of electrical cross talk happening between some MC locking signals and the seismometer channels.

    So we should not trust the signals from the STS seismometers. This needs a further investigation.

(2) We looked at the OSEM and oplev signals of some other suspended optics, and didn't find any corresponding fluctuations.

    The suspensions we checked are ETMX, ETMY, ITMX and MC1.

     None of them showed an obvious sign of the active ground motions in the past 24 hours or so.

Quote from #6266

It seems that somehow the seismic noise became louder from about 1:00 AM.

  6273   Fri Feb 10 15:54:27 2012 steveUpdatePEMAC turned back ON

The air cond was off for 2 hrs.  I just switched it back on at 15:51

  6272   Fri Feb 10 15:52:35 2012 JenneUpdatePEMseismic BLRMS loud too

Quote:

 

 Kiwamu and Steve maybe don't know about how to trend seismic noise. If you just take the mean of the time series, you don't prove that the seismic noise got any higher. The STS has a nominally zero DC output, so the long period level shifts that you see tell you just that there was a DC offset.

This is NOT an increase in seismic noise. To see a seismic trend you should plot the trend of the BLRMS channels that we made especially for this purpose.

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

  6271   Fri Feb 10 15:47:38 2012 ranaUpdatePEMseismic noise back to normal

 

 Kiwamu and Steve maybe don't know about how to trend seismic noise. If you just take the mean of the time series, you don't prove that the seismic noise got any higher. The STS has a nominally zero DC output, so the long period level shifts that you see tell you just that there was a DC offset.

This is NOT an increase in seismic noise. To see a seismic trend you should plot the trend of the BLRMS channels that we made especially for this purpose.

  6270   Fri Feb 10 15:46:59 2012 steveUpdateSUSruby wire standoff

Finally I found a company who can do Koji's improved  -hard to make-  specification on ruby or sapphire wire standoff.

NOT POLISHED excimer laser cut, wire groove radius R 0.0005" + - 0.0002"

$250 ea at 50 pieces of order

  6269   Fri Feb 10 11:46:44 2012 steveUpdateIOOseismic noise back to normal

The shaking has stopped at 9:32am  The AC was turned back on at 11:30am  We still do not have any explanation

 

Attachment 1: seism4h.png
seism4h.png
Attachment 2: seis60s.png
seis60s.png
Attachment 3: oneday.png
oneday.png
ELOG V3.1.3-